OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 2 - 8 + 32 - 128 + ... PLEASE HELP ME

OpenStudy (anonymous):

wait I think I got it but I'm not sure if it would be 2(-4)^n or 2(-4)^n-1

OpenStudy (akashdeepdeb):

The common ratio is -4 right? a = 2 r = -4 n = GOES ON FOREVER starting from 1 :) nth term = ar^(n-1) = 2.-4^(n-1) So summation notation would be giving the sum of all the n terms [One by one] \[\sum_{1}^{n} 2.(-4)^(n-1)\] Understood? :)

OpenStudy (anonymous):

AHHHH OKAY!!! XD that;s why it would be n-1!!! I understand way more now :) thank you!!

OpenStudy (akashdeepdeb):

:)